I have an area 128 feet x 168 feet.
I would like to fill this area with sand to a depth of 3 inches.
How much sand will I need?

Answer by vihits13(20)

this question is basiclly asking you to find the volume.
the formula to use is Base Area * height
Base Area= 128*168 = 21,504 square feet
Height = 3 inches
Now convert 21504 feet^2 into inches.
21504*144= 3096576 square inches.
Volume = 3096576*3= 9,289,728
you will need 9,289,728 cubic inches of sand.